OpenStudy (anonymous):

The sum of two consecutive even integers is 154. Find the integers.

OpenStudy (goldphenoix):

You can use the equation: (x) + (x+2) = 154. Now solve for x. Do you know how to solve this equation?

OpenStudy (goldphenoix):

Yes, no, or are you still solving it?

OpenStudy (anonymous):

well i dont really know how to solve it

OpenStudy (loser66):

what does consecutive mean?

OpenStudy (anonymous):

numbers lik 2,4,6,8

OpenStudy (goldphenoix):

Well, that is even consecutive. Though you get what consecutive mean.

OpenStudy (anonymous):

yea...so how do you solve the equation?

OpenStudy (goldphenoix):

I'm not the best at explaining on how I obtain that equation, but I'm sure Loser66 can.

OpenStudy (loser66):

@Jewell64 what if they are not even, how can I know whether they are consecutive or not,? like 1, 2, 3.....

OpenStudy (loser66):

No, no, @GoldPhenoix I am asking her the things I don't know

OpenStudy (goldphenoix):

To solve the equation, pretend the parenthesis isn't there. The reason why I put parenthesis is to split them up into integer. (x) = Smallest integer (x+2) = Largest integer Add 2 because it says even. ^_^ So your equation would look like: x + x + 2 = 154.

OpenStudy (anonymous):

well. consecutive even are all the even numbers lik 2,4,6,8,10 etc. Consecutive odd are all the odd numbers @Loser66

OpenStudy (goldphenoix):

Yep. Now back to the equation: x + x + 2 = 154 The first step is to combine like term. So x + x = ?

OpenStudy (loser66):

thank you

OpenStudy (goldphenoix):

Also, if you don't understand how to do it, then just say so. :)

OpenStudy (anonymous):

isnt x the same as 1?

OpenStudy (goldphenoix):

Not necessarily. x by itself is 1x. x is a variable. It can be any number. Hmm... You can say x is 1... for now. 1 + 1 =?

OpenStudy (anonymous):

2 @GoldPhenoix

OpenStudy (anonymous):

so would it be 2x+2=154? @GoldPhenoix

OpenStudy (goldphenoix):

Yes! Perfect. You want x by itself in one side of the equation. So you want to get rid of 2. To get rid of 2, you want to do inverse operation. So 154 - 2 =?

OpenStudy (anonymous):

so subtract 2 from both sides and u end up with 2x=152. then divide by 2 right?

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

Yep. What do you get?

OpenStudy (anonymous):

x=76

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

Perfect. Now like I said before: x = Smallest number x + 2 = Largest number. Since 76 is the smallest number, what's the largest number?

OpenStudy (anonymous):

78? @GoldPhenoix

OpenStudy (goldphenoix):

Good job! 76, and 78 are your numbers! To check your work just see: Are those consecutive numbers? Do these two add up to 154? You don't have to reply, just check to see if it's right. ^_^

OpenStudy (anonymous):

they do!! thank you so much! @GoldPhenoix

OpenStudy (goldphenoix):

No problem!~